# Uncovering Grounding IDs: How External Cues Shape Multimodal Binding

### v3

@@ 4.2 Layerwise Analysis of Grounding ID Emergence
Layerwise logit difference. To locate where in the network the swapped grounding takes effect, we apply a logit-lens probe to the response token at each layer $\ell$. Let $L^{(\ell)}(x\mid c^{*})$ be the unnormalized logit for token $x$ decoded from $h^{(\ell)}(c^{*})$. We define the layerwise logit difference
$\Delta L^{(\ell)}\;=\;L^{(\ell)}\!\bigl(\mathbf{o}^{s}_{\sim s}\mid c^{*}\bigr)\;-\;L^{(\ell)}\!\bigl(\mathbf{o}^{\sim s}_{s}\mid c^{*}\bigr),$
averaged over all valid symbol–object pairs. Positive values of $\Delta L^{(\ell)}$ indicate that, at layer $\ell$, the representation favors the grounded object over the local object.
Next, we investigate in which layers the model begins to predict the intervened grounded object ($\textbf{o}^{s}_{\sim s}$) rather than the adjacent object ($\textbf{o}^{\sim s}_{s}$). The goal is to identify the layer at which the model’s prediction shifts toward the object carrying the transferred binding in the patched context. In this experiment, we use a monochrome dataset for simplicity. We apply the logit lens technique (nostalgebraist20_logit_lens) to the single-word response token at each layer $\ell$. Let $L^{(\ell)}(x\mid c^{*})$ be the unnormalized logit for token $x$ decoded from $h^{(\ell)}(c^{*})$. We define the layerwise logit difference
@@
to compute the prediction tendency between $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{\sim s}_{s}$ in each patched context, and then average across all valid symbol–object pairs.
As shown in Fig. 5(a), this difference becomes positive in the later layers (20–27), indicating that the model increasingly favors the grounded object after the intervention. This result complements the representational trend in Fig. 3a, where alignment for structured inputs increases in the same higher layers, showing that the late-layer alignment shifts are accompanied by causal evidence of Grounding ID usage.
We also investigate which attention heads are most responsible for propagating Grounding IDs as opposed to relying on local visual proximity. For each head and layer, we compute the difference in attention from the response token to visual tokens corresponding to $\textbf{o}^{s}_{\sim s}$ versus $\textbf{o}^{\sim s}_{s}$ across samples. To quantify consistency, we divide the mean difference by its standard deviation, yielding a signal-to-noise ratio that reflects how reliably a head prefers grounded objects. To emphasize attention to meaningful regions, this ratio is multiplied by the head’s average attention weight. Let $\alpha^{(\ell,h)}(\mathbf{r}\!\to\!\mathbf{o})$ denote the attention weight from the response token $\mathbf{r}$ to the patch of object $\mathbf{o}$ at layer $\ell$, head $h$. For each head, we compute
$S^{(\ell,h)}\;=\;\frac{\mathbb{E}\!\left[\alpha^{(\ell,h)}(\mathbf{r}\!\to\!\mathbf{o}^{s}_{\sim s})-\alpha^{(\ell,h)}(\mathbf{r}\!\to\!\mathbf{o}^{\sim s}_{s})\right]}{\operatorname{Std}\!\left[\alpha^{(\ell,h)}(\mathbf{r}\!\to\!\mathbf{o}^{s}_{\sim s})-\alpha^{(\ell,h)}(\mathbf{r}\!\to\!\mathbf{o}^{\sim s}_{s})\right]}\cdot\mathbb{E}\!\left[\alpha^{(\ell,h)}(\mathbf{r}\!\to\!\text{image})\right],$ (2)
where expectations are taken over samples and valid symbol–object pairs. Heads with high $S^{(\ell,h)}$ consistently prefer grounded objects over merely co-located ones and are interpreted as key carriers of Grounding IDs. The resulting scores are visualized in Fig. 5(b), showing that Grounding IDs cause significant attention shifts in certain heads, particularly in middle layers (layer 16), which attend more to grounded objects.
@@ 4.3 The Characteristics of Grounding IDs
binding_id shows that, in LLMs, Binding IDs linking entities and their attributes are largely context independent. In contrast, our causal mediation experiment demonstrates that Grounding IDs are directly predictable from their corresponding symbols. To probe this relationship, we analyze the relational structure of symbols and their induced Grounding IDs. Concretely, for each pair of symbols (e.g., &, #), we compute the difference between their symbol patch activations. For the same pairs, we also compute the difference between the corresponding object patch activations. Averaging these differential vectors across the dataset cancels out confounding factors such as shape and position, yielding two structured spaces: one defined by symbol differences and one by Grounding ID differences. Fig. 6(a) reports cosine similarities between these two spaces, revealing strong correspondence between the symbol space and the Grounding ID space.
@@ 5 Behavioral Implications of Grounding IDs
Our observations in Section 3 exhibit a reduced modality gap and increased attention between related partitions in samples with external cues. The causal analysis in Section 4 further reveals the presence of latent identifiers that bind to the embeddings of the corresponding partitions. Together, these results describe a mechanism by which external cues improve visual reasoning: they slightly modify the model’s internal dynamics and equip visual embeddings with auxiliary identifiers that strengthen the model’s grounding ability. To examine how this property affects downstream behavior, we first evaluate the impact of Grounding IDs on hallucination mitigation during long caption generation, since this task directly depends on grounding. We then assess the effect of aligned multimodal cues on broader visual reasoning tasks.
@@ 5.1 Hallucination Mitigation
In Section 3, we reported overall attention enhancement across the bound partitions. To test whether this partition-based effect also helps LVLMs remain focused on the image in long responses, we measure cross-attention during generation. Using the same synthetic dataset as in Section 3, the model is asked to produce detailed descriptions of each image based on the provided external cues. We compute cross-attention from generated tokens to image patches using a sliding window, taking the maximum value within each window of size 5. Plotting these values against token position (Fig. 7) reveals a consistent decline: tokens appearing later in the response attend less to the image, indicating that visual grounding diminishes as generation progresses. Importantly, the structured baseline shows both higher initial attention and a slower rate of decline compared to the baseline. This indicates that external cues sustain grounding over longer spans of text.
@@
# Obj. Method Precision Recall F1 Accuracy
@@
Structured 0.65 0.59 0.62 0.40
Prior work attributes hallucinations in LVLMs to the decay of visual attention and increasing reliance on language priors (m3id; opera). Motivated by this, we evaluate hallucination on the controlled datasets from Section 3, measuring accuracy, precision, recall, and F1 for object descriptions, as reported in Table 1. Structured inputs consistently improve performance across all metrics, with particularly strong gains in precision, which is most relevant for avoiding mentions of nonexistent objects. For images with 10 objects, baseline recall is slightly higher, but as the number of objects increases, structured inputs outperform the baseline across all criteria. The performance gap due to external structure also widens as the number of objects increases. Notably, multimodal cues are considerably more effective than unimodal cues. These results highlight that sustained cross-modal grounding directly improves faithfulness of generated text.
Finally, we evaluate on large-scale hallucination benchmarks using MS-COCO images (ms_coco), with performance assessed by the hallucination-specific metric CHAIR (chair). We conduct experiments on two recent LVLMs, Qwen2.5-VL (qwen2.5_vl), and LLaVA-1.5 (llava1.5). We observe that for natural scenes, simple horizontal lines are insufficient, and grid-based partitions provide more effective structure for sequential scanning of the image. To ensure visibility against diverse backgrounds, we add thin white margins around the scaffolding.
Results (Fig. 8 and Table 2) show substantial reductions in both CHAIRs and CHAIRi compared to baseline. Importantly, unlike most existing techniques for hallucination mitigation, the proposed approach is also applicable to black-box, closed-source models such as GPT-4o (gpt4o) and Gemini-2.5-Pro (gemini2.5), which are considered strong models for visual reasoning. Notably, our simple strategy outperforms or matches specialized hallucination-mitigation methods such as VCD (vcd), OPERA (opera), and SPARC (sparc), while requiring no additional inference modules and maintaining near-zero computational overhead. We also report results on another commonly used benchmark, POPE (pope), in Appendix A.6
@@ 5.2 Visual Reasoning Performance
We evaluate Grounding IDs on two visual reasoning benchmarks related to spatial grounding: counting and visual search. We conduct experiments on the datasets introduced in viser. The counting dataset consists of synthetic 2D scenes with 10 to 20 circles, increasing in steps of 2, with 100 images for each count level (600 total). The visual search dataset contains conjunctive search scenes with 20, 30, 40, and 50 objects. For each difficulty level, 100 images are created: 50 with the target present and 50 absent. Accuracy is computed separately for visible and invisible cases and combined using the harmonic mean.
We evaluate three models: Qwen-3B, Qwen-7B, and GPT-4o. Grounding IDs increase accuracy on both tasks and outperform the VISER baseline and the unstructured input. The gains depend on two conditions: each partition must have a distinct identifier, and the same identifiers must appear in the image and the prompt. The results are shown in Table 3.
@@
Counting Accuracy Visual Search Accuracy
Model Baseline VISER Grounding IDs Baseline VISER Grounding IDs
Qwen2.5-VL (3B) 30.00 37.83 43.00 0.00 37.83 45.96
Qwen2.5-VL (7B) 29.67 43.33 53.00 30.00 40.00 52.25
GPT-4o 10.50 26.50 32.33 49.41 73.40 80.62
6 Related Work
@@
Interpretability research investigates where and how information is represented and routed in neural networks. For LLMs, attention analyses study which tokens heads attend to and how this shapes predictions (chefer21_generic_attention; clark19_bert_attention), while circuit-oriented methods identify causal substructures and enable editing through activation or circuit discovery and factual localization (conmy23_automated_circuit_discovery; meng22_rome_factual_edit; geva23_dissect_recall). Probing tools such as the logit lens decode intermediate states to reveal how token predictions evolve across layers (nostalgebraist20_logit_lens). For VLMs, recent work explores where visual information is stored and how it is transferred into the language pathway (basu24_storage_transfer_vlm), and adapts logit-lens–style spatial probes to study grounding and localization inside multimodal models (jiang24_vl_rep_edit; llava_interp). Overall, the interpretability methods provide the toolset we build on: causal interventions and layerwise probes to expose mechanisms underlying cross-modal binding and grounding.
@@
Recent advances suggest that models internally rely on abstract latent variables to maintain entity–attribute associations. In LLMs, Binding IDs are content-independent identifiers that link entities and attributes through a shared latent code; causal interventions that swap these vectors systematically change the inferred associations (binding_id). Follow-up work localizes a low-rank subspace that causally governs which entity pairs with which attribute (dai24_binding_order). In multimodal models, analogous identifiers appear in VLMs: distinct binding codes are attached to an object’s image tokens and its textual mentions, yielding cross-modal referential consistency (vision_binding_id). Complementary evidence points to symbolic indexing in VLMs, where attention heads compute content-independent spatial indices and later retrieve attributes by these indices, thereby implementing object-centric binding across modalities (visual_symbolic_mechanism). These findings highlight the role of hidden symbolic variables in reasoning. While prior analyses primarily examine existing identifiers, we show that simple input-augmented artifacts can elicit partition-specific identifiers that are linked across modalities and enhance grounding through tighter cross-modal attention and alignment.
@@
Efforts to improve reasoning in LVLMs have followed two main directions. Prompt-focused methods such as chain-of-thought prompting, majority voting, and test-time compute scaling aim to enhance reasoning without altering inputs. However, evaluations on multimodal reasoning benchmarks show that these approaches remain ineffective for tasks requiring visual or spatial reasoning, even when scaled substantially (emma). In contrast, recent studies demonstrate that external scaffolding—adding annotations, grids, or partitions to the input image—consistently improves model performance across counting, spatial relations, and description tasks (forgotten_polygons; viser). While such findings establish the effectiveness of scaffolding, the mechanisms underlying these improvements remain poorly understood.
@@
Hallucination remains a central challenge for LVLMs, mainly arising from the decline of cross-modal attention during long text generation and over-reliance on language priors. Several inference-time methods have been proposed, including VCD (vcd), which down-weights tokens favored under distorted inputs; OPERA (opera), which applies an over-trust penalty and retrospection allocation to rebalance attention; and SPARC (sparc), which progressively recalibrates visual attention to maintain relevance in long captions. Our work complements these approaches by identifying how visual scaffolding induces Grounding IDs that naturally sustain cross-modal binding, thereby reducing hallucination without additional inference tools.
7 Conclusion
In this work, we introduce a conceptual framework to explain how multimodal aligned external cues induce abstract Grounding IDs across related partitions. Through attention and embedding alignment analysis, we showed that these identifiers propagate across modalities, supporting partition-specific grounding and reducing the modality gap. Activation patching interventions confirmed that the associations are mediated by abstract identifiers rather than local features. Empirical evaluations further demonstrated that Grounding IDs enhance cross-modal attention, yield more faithful image descriptions, and reduce hallucinations. Since the approach relies only on simple, content-independent structures, it provides a model-agnostic strategy applicable to a wide range of tasks and models, including closed-source LVLMs.
Appendix A Appendix
A.1 Visual Traversal and Positional Inductive Bias

### v4

@@ 4.2 Layerwise Analysis of Grounding ID Emergence
Layerwise logit difference. Here, we investigate in which layers the model begins to predict the intervened $\mathbf{o}^{s}_{\sim s}$ (bound object to $s$) rather than $\mathbf{o}^{\sim s}_{s}$ (adjacent object to $s$). The goal is to identify the layer at which the model’s prediction shifts toward the object carrying the transferred binding in the patched context. In this experiment, we use a monochrome dataset for simplicity. We apply the logit lens technique (nostalgebraist, 2020) to the single-word response token at each layer $\ell$. Let $L^{(\ell)}(x\mid c^{*})$ be the unnormalized logit for token $x$ decoded from $h^{(\ell)}(c^{*})$. We define the layerwise logit difference
@@
to compute the prediction tendency between $\mathbf{o}^{s}_{\sim s}$ and $\mathbf{o}^{\sim s}_{s}$ in each patched context, and then average across all valid symbol–object pairs. Positive values of $\Delta L^{(\ell)}$ indicate that, at layer $\ell$, the representation favors the bound object over the local object.
As shown in Fig. 5(a), this difference becomes positive in the later layers (20–27), indicating that the model increasingly favors the bound object after the intervention. This result complements the representational trend in Fig. 3a, where alignment for structured inputs increases in the same higher layers, showing that the late-layer alignment shifts are accompanied by causal evidence of Grounding ID usage.
Responsible attention heads. We also investigate which attention heads are most responsible for propagating Grounding IDs as opposed to relying on local visual proximity. For each head and layer, we compute the difference in attention from the response token to visual tokens corresponding to $\mathbf{o}^{s}_{\sim s}$ vs. $\mathbf{o}^{\sim s}_{s}$ (the bound object vs. the adjacent one) across samples. To quantify consistency, we divide the mean difference by its standard deviation, yielding a signal-to-noise ratio that reflects how reliably a head prefers bound objects. To emphasize attention to meaningful regions, this ratio is multiplied by the head’s average attention weight. Let $\alpha^{(\ell,h)}(\mathbf{r}\!\to\!\mathbf{o})$ denote the attention weight of the response token $\mathbf{r}$ on the patch corresponding to object $\mathbf{o}$ at layer $\ell$, head $h$. For each head, we compute
$S^{(\ell,h)}$ $=\frac{\mathbb{E}\!\left[\alpha^{(\ell,h)}(\mathbf{r}\!\to\!\mathbf{o}^{s}_{\sim s})-\alpha^{(\ell,h)}(\mathbf{r}\!\to\!\mathbf{o}^{\sim s}_{s})\right]}{\operatorname{Std}\!\left[\alpha^{(\ell,h)}(\mathbf{r}\!\to\!\mathbf{o}^{s}_{\sim s})-\alpha^{(\ell,h)}(\mathbf{r}\!\to\!\mathbf{o}^{\sim s}_{s})\right]}$ (2)
$\quad\cdot\mathbb{E}\!\left[\alpha^{(\ell,h)}(\mathbf{r}\!\to\!\text{image})\right].$
where expectations are taken over samples and valid symbol–object pairs. Heads with high $S^{(\ell,h)}$ consistently prefer bound objects over merely co-located ones and are interpreted as key carriers of Grounding IDs. The resulting scores are visualized in Fig. 5(b), showing that Grounding IDs cause significant attention shifts in certain heads, particularly in middle layers (layer 16), which attend more to bound objects.
@@ 4.3 The Characteristics of Grounding IDs
Feng and Steinhardt (2023) shows that, in LLMs, Binding IDs linking entities and their attributes are largely context independent. In contrast, our causal mediation experiments demonstrate that Grounding IDs are directly predictable from their corresponding symbols, suggesting a mechanism closer to lexical binding (Gur-Arieh et al., 2025). To probe the characteristics of Grounding IDs, we analyze the relational similarity between symbols and their induced Grounding IDs. Concretely, for each pair of symbols (e.g., &, #), we compute the difference between their symbol patch activations. For the same pairs, we also compute the difference between the corresponding object patch activations. Averaging these differential vectors across the dataset cancels out confounding factors such as shape and position, yielding two structured spaces: one defined by symbol differences and one by Grounding ID differences. Fig. 6(a) reports cosine similarities between these two spaces, revealing a strong correspondence between the symbol space and the Grounding ID space, which further supports a lexical-style binding mechanism.
@@ 5 Behavioral Implications of Grounding IDs
Our observations in Section 3 exhibit a reduced modality gap and increased attention between related partitions in samples with external cues. The causal analysis in Section 4 further reveals the presence of latent identifiers that bind to the embeddings of the corresponding partitions. To examine how this property affects downstream behavior, we first evaluate the impact of Grounding IDs on hallucination mitigation during long caption generation, since this task directly depends on grounding. We then assess the effect of aligned multimodal cues on broader visual reasoning tasks.
@@ 5.1 Hallucination Mitigation
In Section 3, we reported overall attention enhancement across the bound partitions. To test whether this partition-based effect also helps LVLMs remain focused on the image in long responses, we measure cross-attention during generation. Using the same synthetic dataset as in Section 3, the model is asked to produce detailed descriptions of each image based on the provided external cues. We compute cross-attention from generated tokens to image patches using a sliding window, taking the maximum value within each window of size 5. Plotting these values against token position (Fig. 7) reveals a consistent decline: tokens appearing later in the response attend less to the image, indicating that visual grounding diminishes as generation progresses. Importantly, the structured case shows both higher initial attention and a slower rate of decline compared to the baseline. This indicates that external cues sustain grounding over longer spans of text.
@@
# Obj. Method Precision Recall F1 Acc.
@@
Structured (both) 0.65 0.59 0.62 0.40
Prior work attributes hallucinations in LVLMs to the decay of visual attention and increasing reliance on language priors (Favero et al., 2024; Huang et al., 2024). Motivated by this, we evaluate hallucination on the controlled datasets from Section 3, as reported in Table 1. Structured modalities consistently improve performance across all metrics, with particularly strong gains in precision, which is most relevant for avoiding mentions of nonexistent objects. For images with 10 objects, baseline recall is slightly higher, but as the number of objects increases, structured inputs outperform the baseline across all criteria. The performance gap due to external structure also widens as the number of objects increases. Notably, multimodal cues are considerably more effective than unimodal cues. These results highlight that sustained cross-modal grounding directly improves faithfulness of generated text.
Finally, we evaluate on large-scale hallucination benchmarks using MS-COCO images (Lin et al., 2014), with performance assessed by the hallucination-specific metric CHAIR (Rohrbach et al., 2018). We conduct experiments on two recent LVLMs, Qwen2.5-VL (Bai et al., 2025), and LLaVA-1.5 (Liu et al., 2023a). We observe that for natural scenes, simple horizontal lines are insufficient, and grid-based partitions provide more effective structure for sequential scanning of the image. To ensure visibility against diverse backgrounds, we add thin white margins around the scaffolding.
Results (Fig. 8 and Table 2) show substantial reductions in both CHAIRs and CHAIRi compared to baseline. Importantly, unlike most existing techniques for hallucination mitigation, the proposed approach is also applicable to black-box, closed-source models such as GPT-4o (Hurst et al., 2024) and Gemini-2.5-Pro (Comanici et al., 2025), which are considered strong models for visual reasoning. Notably, our simple strategy outperforms or matches specialized hallucination-mitigation methods such as VCD (Leng et al., 2024), OPERA (Huang et al., 2024), and SPARC (Jung et al., 2025), while requiring no additional inference modules and maintaining near-zero computational overhead. We also report results on another commonly used benchmark, POPE (Li et al., 2023), in Appendix G.
@@ 5.2 Visual Reasoning Performance
We evaluate the effectiveness of modality-aligned cues on two visual reasoning benchmarks, using the same experimental setup as Izadi et al. (2025), and compare it with the existing method, VISER. The results for the counting and visual search tasks are shown in Table 3.
We evaluate three models: Qwen-3B, Qwen-7B, and GPT-4o. Grounding IDs increase accuracy on both tasks and outperform the VISER baseline and the unstructured input. The gains depend on two conditions: each partition must have a distinct identifier, and the same identifiers must appear in the image and the prompt.
@@
Counting Accuracy
Model Baseline VISER Grounding IDs
Qwen2.5-VL (3B) 30.00 37.83 43.00
Qwen2.5-VL (7B) 29.67 43.33 53.00
GPT-4o 10.50 26.50 32.33
Visual Search Accuracy
Model Baseline VISER Grounding IDs
Qwen2.5-VL (3B) 0.00 37.83 45.96
Qwen2.5-VL (7B) 30.00 40.00 52.25
GPT-4o 49.41 73.40 80.62
6 Conclusion
In this work, we introduce a conceptual framework to explain how multimodal aligned external cues induce abstract Grounding IDs across related partitions. Through attention and embedding alignment analysis, we showed that these identifiers propagate across modalities, supporting partition-specific grounding and reducing the modality gap. Activation patching interventions confirmed that the associations are mediated by abstract identifiers rather than local features. Empirical evaluations further demonstrated that Grounding IDs enhance cross-modal attention, yield more faithful image descriptions, reduce hallucinations, and improve visual reasoning. Since the approach relies only on simple, content-independent structures, it provides a model-agnostic strategy applicable to a wide range of tasks and models, including closed-source LVLMs.
Impact Statement
This work aims to advance the understanding and interpretability of vision–language models by studying how simple, aligned external cues improve multimodal grounding and reduce hallucinations. The techniques explored are model-agnostic and rely on lightweight input modifications, which may contribute to safer and more reliable deployment of multimodal systems in applications that require faithful visual grounding. We do not anticipate significant negative societal impacts beyond those already associated with large-scale vision–language models, and this work does not introduce new data sources, learning objectives, or deployment scenarios that raise additional ethical concerns.
Appendix A Related Work
@@
Interpretability research investigates where and how information is represented and routed in neural networks. For LLMs, attention analyses study which tokens heads attend to and how this shapes predictions (Chefer et al., 2021; Clark et al., 2019), while circuit-oriented methods identify causal substructures and enable editing through activation or circuit discovery and factual localization (Conmy et al., 2023; Meng et al., 2022; Geva et al., 2023). Probing tools such as the logit lens decode intermediate states to reveal how token predictions evolve across layers (nostalgebraist, 2020). For VLMs, recent work explores where visual information is stored and how it is transferred into the language pathway (Basu et al., 2024), and adapts logit-lens–style spatial probes to study grounding and localization inside multimodal models (Jiang et al., 2024; Neo et al., 2024). Overall, the interpretability methods provide the toolset we build on: causal interventions and layerwise probes to expose mechanisms underlying cross-modal binding and grounding.
@@
Recent advances suggest that models internally rely on abstract latent variables to maintain entity–attribute associations. In LLMs, Binding IDs are content-independent identifiers that link entities and attributes through a shared latent code; causal interventions that swap these vectors systematically change the inferred associations (Feng and Steinhardt, 2023). Follow-up work localizes a low-rank subspace that causally governs which entity pairs with which attribute (Dai et al., 2024). In multimodal models, analogous identifiers appear in VLMs: distinct binding codes are attached to an object’s image tokens and its textual mentions, yielding cross-modal referential consistency (Saravanan et al., 2025). Complementary evidence points to symbolic indexing in VLMs, where attention heads compute content-independent spatial indices and later retrieve attributes by these indices, thereby implementing object-centric binding across modalities (Assouel et al., 2025). These findings highlight the role of hidden symbolic variables in reasoning. While prior analyses primarily examine existing identifiers, we show that simple input-augmented artifacts can elicit partition-specific identifiers that are linked across modalities and enhance grounding through tighter cross-modal attention and alignment.
@@
Efforts to improve reasoning in LVLMs have followed two main directions. Prompt-focused methods such as chain-of-thought prompting, majority voting, and test-time compute scaling aim to enhance reasoning without altering inputs. However, evaluations on multimodal reasoning benchmarks show that these approaches remain ineffective for tasks requiring visual or spatial reasoning, even when scaled substantially (Hao et al., 2025). In contrast, recent studies demonstrate that external scaffolding—adding annotations, grids, or partitions to the input image—consistently improves model performance across counting, spatial relations, and description tasks (Rudman et al., 2025; Izadi et al., 2025). While such findings establish the effectiveness of scaffolding, the mechanisms underlying these improvements remain poorly understood.
@@
Hallucination remains a central challenge for LVLMs, mainly arising from the decline of cross-modal attention during long text generation and over-reliance on language priors. Several inference-time methods have been proposed, including VCD (Leng et al., 2024), which down-weights tokens favored under distorted inputs; OPERA (Huang et al., 2024), which applies an over-trust penalty and retrospection allocation to rebalance attention; and SPARC (Jung et al., 2025), which progressively recalibrates visual attention to maintain relevance in long captions. Our work complements these approaches by identifying how visual scaffolding induces Grounding IDs that naturally sustain cross-modal binding, thereby reducing hallucination without additional inference tools.
Appendix B Visual Traversal and Positional Inductive Bias